# Aleth-NeRF: Illumination Adaptive NeRF with Concealing Field Assumption

### 14 Dec 2023 v1

@@ Methods
We start by training vanilla NeRF (Mildenhall et al. 2020) under low-light conditions as a baseline (first row in Table 1). ① Subsequently, we apply various state-of-the-art enhancement methods to post-process the rendered NeRF results (“NeRF + ” in Table 1). ② Then we pre-process low-light images with enhancement methods and then train NeRF on these enhanced images (“* + NeRF” in Table. 1). Notably, we observe that the “NeRF + *” methods tend to underperform compared to the “* + NeRF” methods, especially in perceptual similarity (LPIPS $\downarrow$). This potentially arise from NeRF’s production of poor-quality images in dark scenes, causing enhancement methods to struggle when applied to the generated dark scene by NeRF. ③ At last, we apply two video enhancement methods MBLLEN (Lv et al. 2018) and LLVE (Zhang et al. 2021a) to pre-process low-light images and then trained with NeRF. ④ Above all, our Aleth-NeRF could gain the best performance on most scenes (see Table. 1), and trained in an end-to-end manner without any pre-processing nor post-processing.
Beyond enhancement methods, RAW-NeRF (Mildenhall et al. 2021) also conducted experiments under low-light conditions, due to their model is trained on RAW data and hard to implement on our dataset, a comparison with RAW-NeRF on their proposed dataset scene is shown in Fig. 5, while NeRF and Aleth-NeRF are trained with 8-bits LDR inputs and RAW-NeRF is trained with 16-bits HDR inputs, our Aleth-NeRF could also gain delight enhancement results while save much GPU memory and training time.
Fig. 6 shows qualitative visualization results in “sofa” and “shrub” scenes, with the comparison of current SOTA image enhancement methods Zero-DCE (Guo et al. 2020), SCI (Ma et al. 2022) and video enhancement method LLVE (Zhang et al. 2021a), we can see that our Aleth-NeRF could gain more vivid enhancement results and closer to ground truth, meanwhile also gain much better in multi-view consistency (see depth maps). More visualization results please refer to supplenmentary.
Table 2: Novel view synthesis performance in over-exposure setting by comparing generated images with ground truth normal-light views. We show the mean value of all 5 scenes.
Method PSNR/ SSIM/ LPIPS Method PSNR/SSIM/LPIPS NeRF + HE 12.72/ 0.553/ 0.588 HE + NeRF 15.45/ 0.647/ 0.558 NeRF + MSEC 12.98/ 0.514/ 0.605 MSEC + NeRF 13.89/ 0.558/ 0.541 NeRF + IAT 11.77/ 0.541/ 0.558 IAT + NeRF 14.31/ 0.536/ 0.507 NeRF 6.74/ 0.285/ 0.867 Aleth-NeRF 15.68/ 0.704/ 0.512
@@
Then we show novel view synthesis results in over-exposure scenes, similar to low-light, we choose 3 exposure correction methods to make comparison, including histogram equalization (HE) (Gonzalez and Woods 2006) and two SOTA deep network methods MSEC (Afifi et al. 2021) and IAT (Cui et al. 2022). As it shown in Table. 2, we use vanilla NeRF trained on over-exposure images to serve as a baseline, and adopt exposure correction methods on generated results (denote as “NeRF + *”). Also adopt exposure correction methods to pre-process over-exposure images and then train NeRF on the processed images (denote as “* + NeRF”). Overall, Aleth-NeRF could gain best results in PSNR and SSIM and second best result in LPIPS. Some visualization results are shown in Fig. 6, our Aleth-NeRF could suppress over-bright meanwhile maintain multi-view consistency.
@@ Conclusion
In this paper, we propose Aleth-NeRF model. Inspired by wisdom of ancient Greeks, we introduce a concept: Concealing Field. After incorporating the concealing fields into NeRF, we can effectively address novel view synthesis tasks in scenes with low-light &\&& over-exposure condition.
One limitation is that Aleth-NeRF should be specifically trained for each scene, which is the same as vanilla NeRF. Besides, Aleth-NeRF may fail in scenes with non-uniform lighting conditions (Wang, Xu, and Lau 2022) or shadow conditions (Qu et al. 2017), we’ll solve these in the future.
Acknowledgement
This work was done during first author internship at Shanghai Artificial Intelligence Laboratory, means the corresponding author. This work was partially supported by JST Moonshot R&\&&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015 and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo. And this work is supported in part by the National Key R&\&&D Program of China (NO. 2022ZD0160100).

### 24 Jan 2024 v3

@@ Methods
We start by training vanilla NeRF (Mildenhall et al. 2020) under low-light conditions as a baseline (first row in Table 2). ① Subsequently, we apply various state-of-the-art enhancement methods to post-process the rendered NeRF results (denote as “NeRF + ” in Table 2). ② Then we pre-process low-light images with enhancement methods and then train NeRF on these enhanced images (denote as “* + NeRF” in Table. 2). Notably, we observe that the “NeRF + *” methods tend to underperform compared to the “* + NeRF” methods in low-light setting, especially in perceptual similarity (LPIPS $\downarrow$). This potentially arise from NeRF’s production of poor-quality images in dark scenes, causing enhancement methods to struggle when applied to the generated dark scene by NeRF. ③ At last, we apply two video enhancement methods MBLLEN (Lv et al. 2018) and LLVE (Zhang et al. 2021a) to pre-process low-light images and then trained with NeRF. ④ Above all, our Aleth-NeRF could gain the best or second best performance on most scenes (see Table. 2), at the same time Aleth-NeRF does not require any training data as prior knowledge, and trained in an end-to-end manner without any pre-processing nor post-processing.
Beyond enhancement methods, RAW-NeRF (Mildenhall et al. 2021) also conducted experiments under low-light conditions, due to their model is trained on RAW data and hard to implement on our dataset, a comparison with RAW-NeRF on their proposed dataset scene is shown in Fig. 5, while NeRF and Aleth-NeRF are trained with 8-bits LDR inputs and RAW-NeRF is trained with 16-bits HDR inputs, our Aleth-NeRF could also gain delight enhancement results while save much data occupy and training time.
Fig. 6 shows qualitative visualization results in LOM dataset’s “sofa” and “shrub” scenes, with the comparison of current SOTA image enhancement methods Zero-DCE (Guo et al. 2020), SCI (Ma et al. 2022) and video enhancement method LLVE (Zhang et al. 2021a), we can see that our Aleth-NeRF could gain more vivid enhancement results and closer to ground truth, meanwhile also gain much better in multi-view consistency (see depth maps). More visualization results please refer to supplementary part.
@@
Then we show novel view synthesis results in over-exposure scenes, similar to low-light scenes, we make comparison with 3 mainstream exposure correction methods, including traditional histogram equalization (HE) method (Gonzalez and Woods 2006) and two current SOTA deep network methods MSEC (Afifi et al. 2021) and IAT (Cui et al. 2022).
As it shown in Table. 3, we use vanilla NeRF trained on over-exposure images to serve as a baseline. ① Then we adopt exposure correction methods to post-process the generated results of NeRF (denote as “NeRF + *”). ② We also adopt exposure correction methods to pre-process over-exposure images and then train NeRF on the processed images (denote as “* + NeRF”). ③ Our proposed Aleth-NeRF method. could gain best results in PSNR and SSIM, but the performance in LPIPS is not very good. Some visualization results are shown in Fig. 6, our Aleth-NeRF could suppress over-bright meanwhile maintain multi-view consistency.
One major reason for the less pronounced LPIPS results on over-exposed images maybe that Aleth-NeRF’s training directly based on degraded overexposed images. Therefore, unsupervised learning faces challenges in compensating for these degraded pieces of information. Aleth-NeRF’s assumption about darkness is grounded in the Concealing Field present in the air, we believe that a more effective modeling of darkness might offer improved solutions for addressing and handling overexposure issues.
@@ Conclusion
In this paper, we propose Aleth-NeRF model. Inspired by wisdom of ancient Greeks, we introduce a concept: Concealing Field. After incorporating the concealing fields into NeRF, we can effectively address novel view synthesis tasks in scenes with low-light &\&& over-exposure conditions. Additionally, we proposed several unsupervised loss functions to constrain the generation of the concealing fields, make Aleth-NeRF could achieve SOTA reults on novel view synthesis under low-light &\&& over-exposure conditions.
One limitation is that Aleth-NeRF should be specifically trained for each scene, which is the same as vanilla NeRF. Besides, Aleth-NeRF may fail in scenes with non-uniform lighting conditions (Wang, Xu, and Lau 2022) or shadow conditions (Qu et al. 2017), we believe that this is also a valuable research topic for future exploration.
Acknowledgements
This work was done during first author internship at Shanghai Artificial Intelligence Laboratory. This work was partially supported by JST Moonshot R&\&&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015 and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo. And this work is supported in part by the National Key R&\&&D Program of China (NO. 2022ZD0160100).
Meanwhile, many thanks to the colleagues from the University of Tokyo MIL Laboratory and Shanghai AI Lab for their active discussions in this work. Additionally, special thanks to Tianhan Xu for his valuable suggestions during the dataset collection phase of this work.